# On the Element-Wise Representation and Reasoning in Zero-Shot Image Recognition: A Systematic Survey

### v2

@@ Abstract
Zero-shot image recognition (ZSIR) aims at empowering models to recognize and reason in unseen domains via learning generalized knowledge from limited data in the seen domain. The gist for ZSIR is to execute element-wise representation and reasoning from the input visual space to the target semantic space, which is a bottom-up modeling paradigm inspired by the process by which humans observe the world, i.e., capturing new concepts by learning and combining the basic components or shared characteristics. In recent years, element-wise learning techniques have seen significant progress in ZSIR as well as widespread application. However, to the best of our knowledge, there remains a lack of a systematic overview of this topic. To enrich the literature and provide a sound basis for its future development, this paper presents a broad review of recent advances in element-wise ZSIR. Concretely, we first attempt to integrate the three basic ZSIR tasks of object recognition, compositional recognition, and foundation model-based open-world recognition into a unified element-wise perspective and provide a detailed taxonomy and analysis of the main research approaches. Then, we collect and summarize some key information and benchmarks, such as detailed technical implementations and common datasets. Finally, we sketch out the wide range of its related applications, discuss vital challenges, and suggest potential future directions.
@@
I Introduction
With the gradual shift from traditional machine learning models to deep neural networks, automated image recognition tasks are gradually achieving accuracy comparable to that of the human brain. The success of deep networks cannot be achieved without two main factors, i.e., massive data along with its annotation and a closed-form environment. The deficiencies are obvious. On the one hand, the cost of image collection is unaffordable; ImageNet [1] contains 14 million images with only 21,814 categories, while most species are much more than that, such as a conservative estimate that there are at least 3 million insect species in the world. On the other hand, image samples of some data are unavailable, such as rare plants, rare medical cases and privacy-protected data. In addition, the requirement for automated models has become more complex now, not only to know what an object is, but also to describe its state [2, 3], e.g., the color of a car, the age of a dog, and so on. The superposition of states and objects means that the cost and difficulty of data collection increase geometrically. As a result, in recent years the interest of researchers has gradually moved from closed-set accuracy to open environment deployment, whose central claim is to endow the model with the ability to discriminate new things without the need to see any visual samples in advance [4, 5, 6]. The way humans think makes the aspiration possible. People can recognize things by extracting and reorganizing their shared characteristics for logical reasoning [7]. For example, we can identify a zebra by combining the morphology of a horse, the stripes of a tiger, and the color of a panda without ever having seen an image of it, and similarly identify a blue car and a red bicycle by combining a red car and a blue bicycle. It implies that human cognition of upper things can be translated into the dismantling and reconstruction of base elements. We have to ask: Can this paradigm of mind be transplanted to the modern automatic recognition system? Naturally, this question motivates the development and application of element-wise learning in open-world recognition tasks.
The objective of element-wise learning is to construct a bottom-up cognitive system by executing element-wise representation and reasoning, i.e., a pyramidal knowledge architecture from micro to macro, attribute to object, component to whole, and local to global. Elements are carriers of properties [8], which describe the meta concepts of a thing, such as materials, colors, patterns, styles, expressions, parts, or functions. The semantic information conveyed by an element varies depending on the specific task and scene. It can be concrete components, e.g., wings and claws, in bird recognition [9] or abstract descriptions, e.g., studying and swimming, in scene classification [10]. Employing element-wise representation to capture discriminative semantic information is already customary for fine-grained classification tasks [11]. And it has been proven effective in many other areas, such as face recognition [12, 13], image retrieval [14] and visual question answering [15]. Element-wise reasoning is the ability to process and apply a complex set of elements. The crux of the extension to the open environment is the combination of different elements to learn new concepts. It is inspired by the hallmark of human intelligence [3], which possesses systematic and structured combinational logic such as predicate logic, boolean logic, and so on [16]. From an early stage, researchers have explored the relationship between combinational elements and objects in visual recognition [17, 18]. Many modern technologies are also designed to follow the idea of combinational reasoning, such as deformable network [19] and graph [7, 20].
Element-wise learning and Zero-Shot Image Recognition (ZSIR) are highly compatible, and the deep integration of the two provides a powerful solution for open-environment deployment. ZSIR employs text as a medium to bridge visual information between known and unknown classes, thus generalizing the model to the unseen domain. Simply put, the shared visual information is compressed into semantic text, and the model unifies the semantic text and image features of the known classes to reason about the unknown classes. As one of the many illuminating solutions, ZSIR has unique advantages. Comparing to other low-shot schemes such as few-shot learning [21] and one-shot learning [22], ZSIR reduces the data requirement to 00 thus completely releasing the data pressure. Comparing to Out-Of-Distribution detection [23], ZSIR assigns precise labels for unknown samples instead of simply categorizing them as unknown. Element-wise learning provides an effective guideline for ZSIR, which profoundly mimics the way humans observe the world, learning the basics through fine-grained language-vision interactions to identify new concepts.
In recent years, ZSIR techniques based on element-wise representation and reasoning have been rapidly developed and widely applied. Unfortunately, the related studies are fragmented and split, lacking a unified and systematic investigation. Although some surveys [4, 24, 6] have been presented, they do not delve in-depth into element-wise techniques. Worse still, they focus only on zero-shot object recognition, which is one of the subtasks of ZSIR, and thus are far from sufficient to show the holistic portrait of ZSIR. To this end, we conduct an extensive survey to thoroughly explore the current situation of the development of element-wise ZSIR techniques. And we attempt to integrate the three mainstream tasks of ZSIR, including object recognition, compositional recognition, and foundation model-based open-world recognition, to provide a more comprehensive view.
Our main contributions are summarized below:
  • We integrate three major zero-shot image recognition tasks, including object recognition, compositional recognition, and foundation model-based open-world recognition, to provide a comprehensive view of element-wise zero-shot learning.
  • We discuss the differences and synergies of several tasks, and analyze the key challenges. We give a detailed introduction to the cutting-edge research techniques and provide a new taxonomy.
  • We collect and present some key information, such as technical details and public datasets.
  • we present the wide range of applications and discuss the potential future directions.
II Description, Comparison, Challenge, and Organization
II-A Description
@@
Figure 1: Three major masks of zero-shot image recognition. Zero-Shot Object Recognition utilizes attributes shared between classes to identify unseen samples. Zero-Shot Compositional Recognition reasons about unseen compositions by learning seen compositions. Foundation model-Based Open-world Recognition exploits the broad fundamentals learned by pre-trained VLMs to implement zero-shot recognition directly in downstream tasks.
Existing surveys only address the topic of zero-shot object recognition, which is far from covering the entire research of ZSIR. In contrast, we integrate three mainstream tasks to provide a more comprehensive view, including object recognition, combination recognition, and foundation model-based open-world recognition. Firstly, we give a detailed introduction to the three tasks.
Zero-Shot Object Recognition (ZSOR). A plain desire is to learn classification in the absence of images to alleviate the dilemma of unavailability of visual data. Humans can recognize 30,000 categories even if some of them have never been seen before [17]. It relies on the full utilization of both verbal and visual media, where we can ”imagine” what an image looks like through its description. Zero-Shot Learning (ZSL) [25] is born to explore the possibility of realizing this capability on automated models. ZSL provides attribute annotations describing the categories, which may be words, sentences, or semantic embeddings. In the training phase, only attribute annotations and images of partial categories are available. The model needs to recognize those categories that have never been seen before in the inference phase. Therefore, the attribute annotations of unseen categories naturally serve as a nexus for conveying the semantics. The setting of ZSL is a bit limited as it only requires testing unseen categories, which is clearly not consistent with real scenarios. Generalized Zero-Shot Learning (GZSL) [4] further extends the setting by requiring both unseen and seen categories to be recognized during the testing period. GZSL is more practical and challenging than ZSL. Another extended setting is Transductive Zero-Shot Learning (TZSL) [26], which relaxes the constraints of ZSL by assuming that images of unseen classes are available but labels are not available during the training phase. The setting of TZSL is too loose to alleviate the problem of data stress, and thus we do not discuss TZSL in this paper. Most studies have accommodated both ZSL and GZSL settings, and thus we collectively refer to them as Zero-Shot Object Recognition (ZSOR). Notably, several surveys on ZSOR have been presented. They comprehensively review the techniques, whereas we focus only on the element-wise studying schemes therein.
Zero-Shot Compositional Recognition (ZSCR). Another fundamental claim is that besides categorizing images, the model can also simply describe relevant states, e.g., whether an apple is ripe or not, or whether a bread is sliced or not. It is an important step for the model to truly understand the world and a key part of building human-like intelligence [27]. States are relevant properties of objects that describe the discrepancy between individuals. It means that while distinguishing between different classes, it is also necessary to recognize the characteristics of individuals in the same class. State recognition can provide more fine-grained information, which is important for model detection, perception and simulation of real environments. However, data collection is a major bottleneck. Each individual has multiple states, which means the number of samples for a complete dataset will grow geometrically. Obviously, it greatly hinders realistic applications. In contrast, humans can reason about the state of an object without having seen all the images. For example, if one has learned about sliced apples, it is easy to recognize sliced oranges. Inspired by that, Compositional Zero-Shot Learning (CZSL) [2, 3, 5] aims to explore the combinational reasoning capabilities of automated models. In the training phase, the model utilizes partial images and corresponding combinational labels for learning. In the testing phase, it needs to recognize different combinations, including seen as well as unseen combinations. Note that in the training set, every state or object has appeared at least once. Conventional CZSL is set up in such a way that the visual samples of unseen combinations are missing but possess a labeling prior, which does not correspond to real scenarios. Therefore, Open-World CZSL [28] extends the setting by expanding the labels tested to the full set, i.e., irrational label combinations, such as sliced cars, may occur, which the model needs to learn to cull out on its own. The technologies of the two are almost crossover, so we collectively refer to them as Zero-Shot Compositional Recognition (ZSCR).
Foundation model-Based Open-world Recognition (FBOR). Despite many remarkable contributions to open deployment, ZSOR and ZSCR are still insufficient to satisfy the notion of complete openness. Their objectives are fixed and the classes to be predicted are well-defined. Can the models learn the broad basics as humans do and thus adapt seamlessly to unknown tasks. The emergence of large vision-language pretraining models (VLMs) [29, 30] offers a promising solution to this problem. During training, the models focus on vision-language interactions and learns large volumes of element-wise knowledge aimlessly, just as humans observe and learn about the world. Surprisingly, the pretrained models have excellent zero-shot generalization ability, which can excel at a variety of unfamiliar image recognition tasks without further training. It has led to a wide range of applications, especially for low-shot problems [31, 32]. However, recently some studies [33, 34] point out that the zero-shot generalization ability of VLMs is not fully exploited. To this end, Foundation model-Based Open-world Recognition (FBOR) is developed to work on finding the limit of the zero-shot performance of foundation models.
Figure 2: An overview of organization.
II-B Comparison
@@
Differences. ZSOR and ZSCR run in relatively closed environments, i.e., the categories or states to be predicted are known. Despite some studies proposing open-world ZSCR, the essence is just to extend the label set from a subset to the full set. FBOR, on the other hand, is task-agnostic, with the need to learn a wider range of knowledge ahead of time to cope with unknown challenges. In general, the element annotations of ZSOR are much larger and more detailed, while ZSCR contains only individual state and category annotations. However, the element annotations of ZSOR are class-level and thus may shift on individual instances, while the annotations of ZSCR are instance-level. In addition, the goal of ZSOR is essentially to recognize objects composed of different attributes, whereas the goal of ZSCR is to recognize combinations composed of different states and objects.
Commonalities. As mentioned earlier, element-wise representation and reasoning is motivation shared by several tasks, i.e., learning the visual meta-concepts corresponding to text annotations, and then combinational reasoning to recognize new upper-level concepts. In addition, many techniques are common across several tasks, such as attention mechanism, graph modeling, etc., which we discuss further in Section VI-C.
II-C Challenge
@@
Fine-grained visual analysis is a critical challenge for Element-Wise Zero-Shot Image Recognition. Element-wise learning requires detailed visual anatomy to echo the text to capture the underlying meta concepts. The ideal solution is a one-to-one visual-text studying strategy, which is impractical to achieve. ZSIR has fine-grained textual information to describe the element-wise semantics. However, constrained by huge labor and time costs, images typically are lacking corresponding element-wise annotations. Therefore, text and visual regions are learned from each other in a weakly supervised manner, and their matching accuracy cannot be guaranteed. In addition, inevitably, irrelevant regions such as the background interfere with the quality of learning. Incorporating fine-grained visual analysis into ZSIR is a promising scenario, which involves visual localization, visual representation, and visual-text correlation modeling. Localization is searching those valuable regions to exclude the interference of noise, forcing the model to focus on the key parts [35, 36]. Representation is processing the contextual relationships between regions to obtain the optimal semantic embeddings [37, 38]. Visual-text correlation modeling is to establish the link between visual and text, involving the design of semantic space and mapping functions [39, 40]. The problems faced include: 1) the complexity of the elements. Compared to class-wise annotations, the number of element-wise annotations tends to be larger and the corresponding visual features are more subtle. For example, in object recognition, the number of attributes can be up to hundreds [9, 41], and they describe semantic information of different dimensions and scales. In addition, the processing of abstract attributes is also a challenge due to the difficulty of finding the corresponding explicit visual regions. 2) visual occlusion and overlap. Visual occlusion occurs between objects and backgrounds or between elements and elements, which may result in the absence of critical visual regions. Visual overlap refers to multiple semantic expressiveness of the same region [42]. For example, when recognizing red car, it is hard to completely divorce the visual information of red and car cause they can be expressed by the same visual part. 3) visual interaction. The independent semantics of visual regions are vague or even ambiguous. For example, if a piece of skin is cut off from a cow’s trunk, it is hard to recognize which part or species it comes from. Thus, the semantic meaning of regions depends to varying degrees on the whole. At the same time, there are structural and physical logical relationships between regions that make independent visual representations suboptimal.
Domain shift is another long-standing issue, which refers to the distributional discrepancy of visual features when mapped into deep semantic space [43]. The conventional paradigm of ZSIR is to map seen visual features to semantic embeddings and associate them to corresponding text annotations. When deployed to the open environment, if unknown visual samples are mapped to the vicinity of similar semantic embeddings, the corresponding text can be indexed and thus the correct label is inferred. However, modern models favor seen classes and thus learns biased mapping functions [44], making the semantic distance between unseen and seen samples too large, i.e., domain shift. It misleads the model to recognize the embeddings of unseen samples as outliers or noise and thus to generate wrong judgments, which is a huge obstacle to generalize the model to the open world. One factor contributing to domain shift is the imaging mechanism and environment [45]. Different cameras, viewpoints, lighting, backgrounds, and many other complex imaging factors may cause the same thing to appear as diverse visual characteristics. For example, the texture of an object photographed in clear weather will be brighter than in dim weather. Humans are able to counteract the style variation easily, but deep models are sensitive. To mitigate this problem, domain adaptation and domain generalization have proposed many fruitful solutions [46, 45, 47]. However, for ZSIR, it is more challenging since domain shift and unknown class prediction need to be accommodated at the same time. Another factor that leads to domain shift is even more troublesome, namely context dependency. Things expressing the same semantic meaning may present very different visual characterizations in different contexts [48, 49]. To give a few examples, the feathers of a chicken are a completely different color when young and when adult; the tail appears hairy and straight on a horse, but smooth and curly on a pig; old presents itself as a rusty patina of metal on a car, but as a fold of skin on a person. This strong contextual dependency is a major hindrance when generalizing a property from one thing to another.
II-D Organization
Fig. 2 illustrates the organization of this paper. Section III provides a comprehensive description of the techniques related to zero-shot object recognition, including the problem formulation and the introduction to the algorithms in the three main directions. Section IV provides a review of the zero-shot compositional recognition approach, also including problem formulation and the introduction to the algorithms in the three main directions. Section V provides an introduction to the background of foundation model-based open-world recognition, followed by a demonstration of three VLM-related zero-shot tasks and recent progress. In section VI, we show the technical details, including commonly used datasets, technical details, and discuss the techniques they share. Section VII shows the related application scenarios. Finally, we share our views on the future directions in section VIII.
III Zero-Shot Object Recognition
III-A Problem Formulation

### v3

@@ Abstract
[…]
@@
1 Introduction
@@
Figure 1: Three tasks of ZSIR: (a) ZSOR utilizes shared attributes/texts to identify unseen categories; (b) ZSCR infers unseen compositions by learning from seen objects and states; and (c) FBOR exploits the broad fundamentals learned by pre-trained VLMs to implement zero-shot recognition directly in downstream tasks.
With the continuous advancement of deep learning techniques, automated image recognition models are gradually becoming comparable to or even better than humans. The success relies on ingenious network architectures and, more importantly, massive data along with accurate annotations. However, the deficiencies are also obvious. On the one hand, the cost of data collection can be huge. For example, despite being a representative large-scale dataset, ImageNet [1] contains 14 million images with 21,814 categories, whose coverage is still far fewer than real-world scenarios, i.e., a conservative estimate suggests that there are at least 3 million species only for the board ‘Insect’ category. On the other hand, image samples of some categories, such as rare/extinct species, rare medical cases, and privacy-sensitive data, are usually limited or even unavailable. Moreover, the requirement for modern deep models has also become more complex, not only to know what an object is but also to describe its accurate states [2, 3], e.g. the ‘Sliced’ apple and an ‘Old’ dog, etc. The superposition of states and objects incurs a geometrical increase in the difficulty of data collection and annotation. In recent years, more and more interests have moved from coarse- to fine-grained scenarios to endow models with the ability to discriminate microscopic characteristics that mimic the human recognition process [4, 5]. For example, one can identify ‘Zebra’ by recognizing an animal with ‘black-and-white’ (color) and ‘stripes’ (physical trait), implying that human cognition of upper concepts can be converted into the dismantling and reconstruction of base elements and further motivate the development of element-wise learning (EWL) in image recognition tasks. Concretely, the objective of EWL is to construct a bottom-up paradigm by executing element-wise representation and reasoning, i.e., a pyramidal knowledge architecture from micro to macro, attribute to object, and local to global. In particular, elements are carriers of properties [6], which describe the meta concepts of samples. The concepts or descriptors conveyed by an element vary depending on the specific tasks, like concrete components, e.g., ‘Wings’ in recognizing bird species [7], or abstract descriptions, e.g., ‘Studying in recognizing scenes [8].
Employing element-wise representation to capture discriminative elements is already customary for fine-grained classification task [9, 4, 5] and has also been proven effective in many other areas, such as face recognition [10], image retrieval [11], and visual question answering [12]. On top of that, element-wise reasoning is employed to assemble a complex set of elements to achieve the recognition. The very early stage began with systematic and structured combinational logic such as predicate logic, boolean logic, and so on [13], and then researchers explored the relationship between combinational elements and objects in visual recognition tasks [14, 15]. Afterward, modern technologies are also designed to follow the idea of combinational reasoning, such as deformable network [16] and graph [17].
Zero-shot image recognition (ZSIR), as a well-defined visual-to-semantics mapping problem adapted for unseen domains, has attracted increasing attention in most recent years and made profound progress [18]. ZSIR employs texts/words as the medium to bridge the image visual disparity between seen and unseen categories, thus generalizing the models to recognize unseen domains. Simply put, the visual features are compressed into the shared semantics, and the model unifies the shared semantics and visual features of the seen categories to reason images from unseen categories. Compared to other low-shot schemes such as few-shot learning [19] and one-shot learning [20], ZSIR reduces the data requirement to zero, thus completely releasing the data collection and annotation pressure. Moreover, compared to another similar open-world problem, i.e., out-of-distribution (OOD) detection [21], ZSIR is designed to assign precise labels for unseen data instead of simply categorizing them as OOD samples, which is usually more favorable in real-world applications.
In practice, ZSIR and EWL are highly compatible in the recognition process that well-mimics the way humans observe the world, i.e., learning the basics through fine-grained visual-to-semantics interaction to identify concepts. Moreover, ZSIR can also be inherently considered as a natural extension of EWL towards the open-environment scenarios, i.e., infer unseen categories with novel combinations of learned base elements. Therefore, their integration can provide a powerful and promising solution for real open-environment deployment. However, existing studies on ZSIR and EWL are usually fragmented and independent of each other, lacking a systematic investigation that summarizes their integration with a unified paradigm. Although some related reviews [22, 23, 24] have been presented for ZSIR in recent years, they do not delve in-depth into element-wise techniques. Worse still, existing reviews focus only on zero-shot object recognition, which is only one subtask of current ZSIR, and thus are constrained from sufficiently presenting the holistic portrait of ZSIR. To fill the gap, we conduct an extensive survey to thoroughly investigate recent advances in the development of element-wise ZSIR approaches and attempt to integrate the three mainstream tasks of ZSIR, including object recognition, compositional recognition, and foundation model-based open-world recognition, to provide a unified and more comprehensive review.
In summary, our contributions are four-fold:
  • We, for the first time, integrate three mainstream ZSIR tasks, i.e., object recognition, compositional recognition, and foundation model-based open-world recognition, to provide a comprehensive review from an element-wise perspective.
  • We discuss the differences and synergies of different ZSIR tasks and analyze the key challenges. We summarize related cutting-edge techniques and provide a new taxonomy of element-wise ZSIR.
  • We collect and summarize some benchmarks, including implementations, datasets, and some more details, as a library for follow-up research.
  • We present the widespread applications, discuss vital challenges and suggest potential future directions.
2 Overview
2.1 Mainstream Tasks
2.1.1 Zero-Shot Object Recognition
Classic zero-shot learning (ZSL) is the most common setting for ZSIR [18], addressing the challenge of classifying images that come from novel unseen categories without observing any prior visual samples before [18]. In ZSL, the annotations such as attributes, words/sentences, or semantic embeddings (i.e., we use semantics uniformly later) are shared to describe categories at a high level. During training, only image samples of seen categories are available, while the model is trained to recognize unseen categories. Hence, the shared semantics serve as the crucial link between seen and unseen categories. To fully adapt to real-world scenarios, generalized ZSL (GZSL) [22] further extends the above setting to a more realistic one by requiring both seen and unseen categories to be jointly recognized during inference. Most recent studies encompass both ZSL and GZSL settings and are collectively referred to as zero-shot object recognition (ZSOR) in this survey (i.e., top-left of Fig. 1).
2.1.2 Zero-Shot Compositional Recognition
Another critical capability in the recent ZSL research is to describe the sample’s relevant states, such as recognizing whether an apple is sliced or raw instead of simply classifying it as an object. Such capability is essential for developing models with a deeper understanding of the world and is also a key component in building human-like intelligence [25]. Inspired by it, compositional learning has emerged in ZSL to explore combinational reasoning capabilities [2, 3, 26]. Concretely, the model is trained from image samples and their corresponding compositional states and objects, and is expected to recognize both seen and novel unseen compositions during inference. In particular, each state and object appears at least once during training. The classic setting assumes that the labels’ prior is available, i.e., some state-object compositions are filtered out, speeding up the inference. In contrast, the open-world setting [27] extends compositional ZSL by expanding the inference scope to the full set, i.e., every state-object composition counts. Therefore, many irrational combinations can occur, and the model is expected to cull out on its own. We collectively refer to them as zero-shot compositional recognition (ZSCR) in this survey (i.e., top-right of Fig. 1).
2.1.3 Foundation Model-Based Open-World Recognition
Both ZSOR and ZSCR fall short of achieving complete openness when faced with the open-world scenario, as their objectives and predictions remain fixed and well-defined. A critical question arises: can models learn broad and foundational knowledge and adapt seamlessly to unknown tasks? The advent of large vision-language pre-trained models (VLMs) [28, 29] offers a promising solution. These models focus on broad vision-language interactions, learning vast amounts of elemental knowledge without a specific goal, i.e., much like humans learn about the world through observation. Remarkably, they exhibit excellent zero-shot generalization ability, performing well on various unseen domain image recognition tasks without further training. This capability has found wide applications, particularly in low-shot scenarios [30, 31]. However, recent studies [32, 33] suggest that the zero-shot potential of VLMs is not fully realized. In response, foundation model-based open-world recognition (FBOR) has been developed to push such a limit in foundation models (i.e., bottom of Fig. 1).
2.2 Comparison
@@
i) Differences. ZSOR and ZSCR both run in relatively narrow environments, i.e., the categories to be predicted are known. Despite some studies proposing open-world ZSCR, the essence is only to extend the label set from a subset to the full set. FBOR, on the other hand, is task-agnostic, with the need to learn a wider range of knowledge ahead of time to cope with unknown challenges. In general, the element annotations of ZSOR are usually larger and more detailed, while ZSCR contains only individual state and object (category) annotations. On the other hand, the element annotations of ZSOR are category-level and thus may shift on individual instances, while the annotations of ZSCR are instance-level. Moreover, the goal of ZSOR is to recognize objects composed of different attributes, whereas the goal of ZSCR is to recognize combinations composed of different states and objects. ii) Commonalities. As mentioned earlier, element-wise representation and reasoning are motivation shared by several tasks, i.e., learning the visual meta-concepts corresponding to text annotations, and then conducting combinational reasoning to recognize novel upper-level concepts. In addition, many techniques are common across several tasks, such as attention mechanism, graph modeling, etc., which we will discuss further in Sect. 6.3.
2.3 Challenge
2.3.1 Fine-Grained Visual Analysis
@@
Element-wise ZSIR requires detailed visual anatomy to echo the semantics and capture underlying meta concepts. ZSIR has fine-grained semantic information to describe the element-wise characteristics. However, due to huge costs, images in hand typically lack enough element-wise annotations. Therefore, semantics and visual local clues are usually learned in a weakly supervised manner, and their matching accuracy cannot be fully guaranteed. Worse still, irrelevant clues, such as the backgrounds, can inevitably interfere with the learning quality. To address these issues, recent studies incorporated fine-grained visual analysis into ZSIR and have made certain progress, particularly in accuracy improvement. Studies of such kind mainly involve visual localization [34, 35], visual representation [36, 37], and visual-semantics correlation modeling [38, 39]. However, based on our observation, three challenges still hinder the fine-grained visual analysis in ZSIR, including i) Elements Complexity, i.e., the number of element-wise annotations tends to be larger compared to category-wise annotations, and the corresponding visual features are much more subtle [7, 40]. In addition, the processing of abstract attributes is also a challenge due to the difficulty of finding the corresponding explicit visual clues; ii) Visual Occlusion and Overlap, i.e., visual occlusion occurs between objects and backgrounds or between elements and elements, which may result in the absence of critical visual clues. Besides, visual overlap refers to multiple semantic expressiveness of the same visual clues [41]. For example, the model can hardly completely decouple the visual clues of ‘red’ and ‘car’ when recognizing a red car; and ii) Visual Interaction, i.e., the independent semantics of visual clues can be vague or even ambiguous. For example, if a piece of skin is cut off from a cow’s trunk, it is hard to recognize which part or even which species the skin comes from. Thus, the semantic meaning of visual clues can also depend on the whole to varying degrees.
2.3.2 Domain Sift
Domain shift is a long-standing challenge in ZSIR, which refers to the distribution bias (i.e., towards seen categories) of visual features when mapped into semantic space [42]. The inducement of domain shift is the absence of visual prior in unseen categories, which makes the models favor too much on seen categories and thus learn biased mapping functions [43]. It further misleads the model in recognizing the embeddings of samples of unseen categories as somewhat seen categories that hinder the model’s generalization ability. Meanwhile, some other factors can also aggravate the domain shift problem. First, different imaging mechanisms cause variational views, lightings, backgrounds, etc. [44], and thus, the visual characteristics can be diverse even for samples of the same category. Domain adaptation and generalization [45, 44, 46] partially mitigates this issue in general recognition tasks, while ZSIR is much more challenging since domain shift and unknown prediction need to be accommodated at the same time. Second and more troublesomely, context dependency, i.e., objects or states expressing the same semantic information may present very different visual characterizations in different contexts [47, 48]. For example, the object ‘tail’ appears hairy and straight on a horse but smooth and curly on a pig, and the state ‘old’ presents itself as the metallic luster of rust on a car, while as the wrinkled skin on a person.
Figure 2: An overview of the organization and taxonomy of this survey.
2.4 Organization
To sum up, Fig. 2 outlines the overall organization as well as the taxonomy of this survey. Concretely, Sect. 3 first covers ZSOR techniques, including problem formulation and key algorithms. Similarly, Sect. 4 reviews ZSCR, and Sect. 5 introduces FBOR. Then, Sect. 6 details technical implementations, standardized datasets, and shared techniques. Sect. 7 explores widespread applications, and Sect. 8 suggests potential future directions.
3 Zero-Shot Object Recognition
3.1 Problem Formulation